OpenStudy (anonymous):

Find the LCM of y^2 - 49 and y^2 + 9y + 14

OpenStudy (anonymous):

I cam up with (y+7)(y-7)(y+7) but that isn't right.

OpenStudy (ybarrap):

y^2 - 49 = y^2 - 7^2 = (y + 7)(y - 7) y^2 + 9y + 14 = (y + 7)(y + 2) LCM = (y^2 - 49)*(y^2 + 9y + 14)/(y + 7) = (y + 7)(y - 7) (y + 7)(y + 2)/(y + 7) LCM = (y + 7)(y - 7)(y + 2)

jimthompson5910 (jim_thompson5910):

y^2 - 49 = (y+7)(y-7) y^2 + 9y + 14 = (y+7)(y+2) collect unique factors and combine them to get the LCM of (y+7)(y-7)(y+2)

OpenStudy (anonymous):

Dang, I almost put that! thanks again. I was close, but no cigar.